A 1000kg rocket in space applies 150000000 n of force in the negative direction. A) find the acceleration of the rocket. B) identify the direction of the acceleration
1
Expert's answer
2020-11-09T06:50:53-0500

As per the given question,

Mass of the rocket "(m)=1000kg"

Force applied by the rockets "(F)=150000000 N"

gravitational acceleration acting on the rocket "(g) =9.8 m\/s^2"

Let the resultant acceleration of the rocket is a, and weight of the rocket is working vertically downwards,

a) "ma=F-mg"

Substituting the values,

"a=\frac{F-mg}{m}"

"\Rightarrow a = \frac{150000000-1000\times 9.8}{1000}m/s^2"

"\Rightarrow a = \frac{149990200}{1000}m/s^2"

"\Rightarrow a=149990.2 m/s^2"

direction of the acceleration will be oppsite to the gravity, means in positive direction.